OpenStudy (anonymous):

Two rooms in your house differ in temperature by 3.0 degrees Fahrenheit. a) What's the temperature difference in Celsius?b)What's the temperature difference in kelvin?

OpenStudy (anonymous):

\[F1=1.8*C1+32\] \[F2=1.8*C2+32\]\[\Delta F=1.8* \Delta C \rightarrow \Delta C= \Delta F /1.8\]

OpenStudy (anonymous):

\[\Delta K= \Delta C\]

OpenStudy (anonymous):

so \[\Delta C=1,67\] and \[\Delta K =3\]